OpenStudy (anonymous):

Find the distance driven, if you drive at 60 mph for 3 1/2 hours. Round to the nearest whole mile. A. 17 mi B. 180 mi C. 210 mi D. 215 mi

OpenStudy (aroub):

d=r*t d=60*7/2 d=210

hero (hero):

Of course, you could also have converted 3 1/2 to 3.5.... d = r*t d = 60 * 3.5 d = 210

hero (hero):

@aroub

OpenStudy (aroub):

I hate dealing with decimals :)

OpenStudy (anonymous):

Thanks :D

hero (hero):

Sometimes, It's easier to deal with them, sometimes it's not. Depends on the situation.